# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.1 Lie dialgebra framework
Using the relation (3.1) valid for any $f\in\mathcal{F}_{D}^{\Gamma}$ and $g\in\Omega_{D}^{\Gamma}$, we we may rewrite the first term on the left hand side as $T$ times a sum over the residues at the points $\zeta_{r}$ for $r\in S$ so that
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda))d\lambda=0\,.$ (3.43)
@@
By definition of the bilinear pairing from Proposition 3.1 we thus have $\langle Y,X\rangle=0$, as desired.
@@
Let us now prove the converse. Namely, let $Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ be arbitrary and suppose that $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$. We must show that, in fact, $Y(\lambda)\in\iota_{\lambda}\Omega_{D}^{\Gamma}$. It follows from Proposition 3.2 that $Y=(Y_{0},Y_{1},\ldots,Y_{N},Y_{\infty})\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ has a unique decomposition as a direct sum of the tuples $\mathcal{P}_{+}(Y)=(\mathcal{P}_{0+}(Y_{0}),\mathcal{P}_{1+}(Y_{1}),\ldots,$ $\mathcal{P}_{N+}(Y_{N}),\mathcal{P}_{\infty+}(Y_{\infty}))\in\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $\mathcal{P}_{-}(Y)=\iota_{\lambda}g$ for some $g\in\Omega_{D}^{\Gamma}$. Now, let $X=\iota_{\lambda}f$ for some $f\in\mathcal{F}_{D}^{\Gamma}$. Since $\langle Y,X\rangle=0$, writing this out explicitly means
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda))d\lambda$
$+T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda))d\lambda=0\,.$ (3.44)
@@
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda))d\lambda=0\,.$ (3.45)
One can then show, along the same lines as the argument in the proof of Proposition 3.1, that if $\mathcal{P}_{+}(Y)\neq 0$ then by picking a suitable $f\in\mathcal{F}_{D}^{\Gamma}$ adapted to $\mathcal{P}_{+}(Y)\in\widetilde{\mathfrak{g}}_{D+}^{(1)}$, one can ensure that the expression on the left hand side is non-zero, which is a contradiction. Therefore, we conclude that $\mathcal{P}_{+}(Y)=0$, and hence $Y\in\iota_{\lambda}\Omega_{D}^{\Gamma}$, as required. The proof of the ‘moreover’ part is completely analogous. ∎
The upshot of Proposition 3.4 is that we are now exactly in the setting recalled in points 1 and 2 of Section 2. In particular, the coadjoint representation $\operatorname{Ad}^{R\ast}:\mathbb{G}_{R}\times V\to V$ of the group $\mathbb{G}_{R}$ on our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$ is given by formula (2.5). This will be used below in Lemma 3.5 to give a concise description of the desired coadjoint orbit for the Lax matrix of the cyclotomic Gaudin model.
@@
$\varphi_{+}=\left(\varphi_{0+},\varphi_{1+},\ldots,\varphi_{N+},\varphi_{\infty+}\right),$ (3.46)
@@
$\begin{split}\varphi_{r+}(\lambda_{r})&=\sum_{n=0}^{\infty}\phi_{r}^{(n)}\lambda_{r}^{n},\quad r\neq\infty\,,\\ \varphi_{\infty+}(\lambda_{\infty})&=\bm{1}+\sum_{n=1}^{\infty}\phi_{\infty}^{(n)}\lambda_{\infty}^{n}\,.\end{split}$ (3.47)
The Lie group $\mathbb{G}_{+}=\widetilde{G}_{D+}^{(0)}$ has a natural action on $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ given by conjugation. This defines the coadjoint representation $\operatorname{Ad}^{\ast}:\mathbb{G}_{+}\times V\to V$ from Section 2 explicitly as
$(\varphi_{+},Y)\longmapsto\operatorname{Ad}^{\ast}_{\varphi_{+}}Y=\left(\varphi_{0+}Y_{0}\varphi_{0+}^{-1},\varphi_{1+}Y_{1}\varphi_{1+}^{-1},\ldots,\varphi_{N+}Y_{N}\varphi_{N+}^{-1},\varphi_{\infty+}Y_{\infty}\varphi_{\infty+}^{-1}\right)\,.$ (3.48)
@@
The orbit of the coadjoint action of $\mathbb{G}_{R}=(\widetilde{G}_{D}^{(0)})_{R}$ on an element $\iota_{\lambda}\Lambda\in V_{-}=\iota_{\lambda}\Omega_{D}^{\Gamma}$ has the explicit form
$\mathcal{O}_{\Lambda}=\big{\{}\mathcal{P}_{-}(\operatorname{Ad}_{\varphi_{+}}^{\ast}\iota_{\lambda}\Lambda);\,\varphi_{+}\in\widetilde{G}_{D+}^{(0)}\big{\}}\,.$ (3.49)
@@
By definition, the coadjoint orbit of $(\widetilde{G}_{D}^{(0)})_{R}$ on any $\iota_{\lambda}\Lambda\in\iota_{\lambda}\Omega_{D}^{\Gamma}$ is given by
$\mathcal{O}_{\Lambda}=\big{\{}\operatorname{Ad}_{\varphi}^{R\ast}\iota_{\lambda}\Lambda;\,\varphi\in(\widetilde{G}_{D}^{(0)})_{R}\big{\}}\,.$ (3.50)
Using the explicit form (2.5) for the coadjoint action of $\mathbb{G}_{R}$ we have
$\operatorname{Ad}^{R\ast}_{\varphi}\iota_{\lambda}\Lambda=\mathcal{P}_{-}(\operatorname{Ad}^{\ast}_{\varphi_{+}}\iota_{\lambda}\Lambda)+\mathcal{P}_{+}(\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda)=\mathcal{P}_{-}(\operatorname{Ad}^{\ast}_{\varphi_{+}}\iota_{\lambda}\Lambda)$ (3.51)
where the last equality follows from the fact that $\iota_{\lambda}\Lambda\in V_{-}=\iota_{\lambda}\Omega_{D}^{\Gamma}$ so that $\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda\in V_{-}$ also and hence $\mathcal{P}_{+}(\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda)=0$. ∎
@@
In the remaining sections we will put this setup to use to describe the Lax matrix of the cyclotomic Gaudin model (1.16) as a point in a coadjoint orbit $\mathcal{O}_{\Lambda}$ for some suitable $\Lambda\in\Omega_{D}^{\Gamma}$ and then introduce the ingredients from point 3 in Section 2 to derive the associated Lax equations.
@@ 3.2 Lax matrix
Our algebraic setup covers the case of the cyclotomic Gaudin model with arbitrary multiplicities. However, as mentioned at the start of this section, from now on we will restrict to the case with simple poles at all $\omega^{k}\zeta_{r}$, $k\in\{0,\ldots,T-1\}$, $r\in\{1,\ldots,N\}$, and double poles at the origin and at infinity, since this is the setting required for our examples in Section 4. The discussion in the remaining sections is easily generalised to the case of arbitrary multiplicities.
To describe a coadjoint orbit $\mathcal{O}_{\Lambda}\in\iota_{\lambda}\Omega_{D}^{\Gamma}$ from Lemma 3.5 corresponding to the Lax matrix (1.16) of the cyclotomic Gaudin model, we fix a non-dynamical element $\Lambda\in\Omega_{D}^{\Gamma}$ with the same pole structure as (1.16), namely we introduce
$\Lambda(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{N}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{r}}{\lambda-\omega^{k}\zeta_{r}}+\Lambda_{\infty}\in\Omega_{D}^{\Gamma}\,.$ (3.53)
@@
$\mathcal{P}_{-}\left(\varphi_{+}\,\iota_{\lambda}\left(\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{N}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{r}}{\lambda-\omega^{k}\zeta_{r}}+\Lambda_{\infty}\right)\,\varphi_{+}^{-1}\right)$
$=\iota_{\lambda}\circ\pi_{\lambda}^{(1)}\left(\frac{A_{0}^{(0)}}{\lambda}+\frac{A_{0}^{(1)}}{\lambda^{2}},\frac{1}{T}\frac{A_{1}}{\lambda-\zeta_{1}},\ldots,\frac{1}{T}\frac{A_{N}}{\lambda-\zeta_{N}},A_{\infty}\right)$
$=\iota_{\lambda}\left(\frac{A_{0}^{(0)}}{\lambda}+\frac{A_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{T-1}\sum_{k=0}^{T-1}\frac{\sigma^{k}A_{r}}{\lambda-\omega^{k}\zeta_{r}}+A_{\infty}\right)$
$\equiv\iota_{\lambda}L(\lambda),$ (3.54)
@@
$\begin{split}&A_{0}^{(0)}=\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1}+\big{[}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1}\big{]}\,,\\ &A_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1}\,,\\ &A_{r}=\phi_{r}\Lambda_{r}\phi_{r}^{-1},\quad r\in\{1,\ldots,N\}\,,\\ &A_{\infty}=\Lambda_{\infty}\,,\end{split}$ (3.55)
with $\phi_{r}^{(0)}$ denoted by $\phi_{r}$, for $r\in\{1,\ldots,N\}$, to simplify notations. Notice that we have the relations $\sigma\phi_{0}^{(0)}=\phi_{0}^{(0)}$ and $\sigma\phi_{0}^{(1)}=\omega\phi_{0}^{(1)}$ for the field elements which ensure that $A_{0}^{(0)}\in\mathfrak{g}^{(0)}$ and $A_{0}^{(1)}\in\mathfrak{g}^{(-1)}$. This gives us the desired parameterisation of the cyclotomic Gaudin Lax matrix

### v2

@@ 3.1 Lie dialgebra framework
Using the relation (3.1) valid for any $f\in\mathcal{F}_{D^{\prime}}^{\Gamma}$ and $g\in\Omega_{D^{\prime}}^{\Gamma}$, we may rewrite the first term on the left-hand side as $T$ times a sum over the residues at the points $\zeta_{r}$ for $r\in S$, so that
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda=0.$ (3.22)
@@
By definition of the bilinear pairing from Proposition 3.1, we thus have $\langle Y,X\rangle=0$, as desired.
@@
Let us now prove the converse. Namely, let $Y\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ be arbitrary and suppose that $\langle Y,X\rangle=0$ for all $X\in\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$. We must show that, in fact, $Y(\lambda)\in\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$. It follows from Proposition 3.2 that $Y=(Y_{0},Y_{1},\dots,Y_{N},Y_{\infty})\in\widetilde{\mathfrak{g}}_{D}^{(1)}$ has a unique decomposition as a direct sum of the tuples $\mathcal{P}_{+}(Y)=(\mathcal{P}_{0+}(Y_{0}),\mathcal{P}_{1+}(Y_{1}),\dots,\mathcal{P}_{N+}(Y_{N}),\mathcal{P}_{\infty+}(Y_{\infty}))\in\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $\mathcal{P}_{-}(Y)=\iota_{\lambda}g$ for some ${g\in\Omega_{D^{\prime}}^{\Gamma}}$. Now, let $X=\iota_{\lambda}f$ for some $f\in\mathcal{F}_{D^{\prime}}^{\Gamma}$. Since $\langle Y,X\rangle=0$, writing this out explicitly means
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda$
$\qquad{}+T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\iota_{\lambda_{r}}g(\lambda)\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda=0.$
@@
$T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda+\!\!\sum_{r\in\{0,\infty\}}\!\!\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}\bigl{(}\mathcal{P}_{r+}(Y_{r}(\lambda_{r}))\iota_{\lambda_{r}}f(\lambda)\bigr{)}{\rm d}\lambda=0.$
One can then show, along the same lines as the argument in the proof of Proposition 3.1, that if $\mathcal{P}_{+}(Y)\neq 0$ then by picking a suitable $f\in\mathcal{F}_{D^{\prime}}^{\Gamma}$ adapted to $\mathcal{P}_{+}(Y)\in\widetilde{\mathfrak{g}}_{D+}^{(1)}$, one can ensure that the expression on the left-hand side is non-zero, which is a contradiction. Therefore, we conclude that $\mathcal{P}_{+}(Y)=0$, and hence $Y\in\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$, as required. The proof of the “moreover” part is completely analogous. ∎
The upshot of Proposition 3.4 is that we are now exactly in the setting recalled in points 1 and 2 of Section 2. In particular, the coadjoint representation $\operatorname{Ad}^{R\ast}\colon\mathbb{G}_{R}\times V\to V$ of the group $\mathbb{G}_{R}$ on our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$ is given by formula (2.5). This will be used below in Lemma 3.5 to give a concise description of the desired coadjoint orbit for the Lax matrix of the cyclotomic Gaudin model.
@@
$\varphi_{+}=(\varphi_{0+},\varphi_{1+},\dots,\varphi_{N+},\varphi_{\infty+}),$ (3.23)
@@
$\varphi_{r+}(\lambda_{r})=\sum_{n=0}^{\infty}\phi_{r}^{(n)}\lambda_{r}^{n},\qquad r\neq\infty,$
$\varphi_{\infty+}(\lambda_{\infty})=\bm{1}+\sum_{n=1}^{\infty}\phi_{\infty}^{(n)}\lambda_{\infty}^{n}.$ (3.24)
The Lie group $\mathbb{G}_{+}=\widetilde{G}_{D+}^{(0)}$ has a natural action on $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ given by conjugation. This defines the coadjoint representation $\operatorname{Ad}^{\ast}\colon\mathbb{G}_{+}\times V\to V$ from Section 2 explicitly as
$(\varphi_{+},Y)\longmapsto\operatorname{Ad}^{\ast}_{\varphi_{+}}Y=\bigl{(}\varphi_{0+}Y_{0}\varphi_{0+}^{-1},\varphi_{1+}Y_{1}\varphi_{1+}^{-1},\dots,\varphi_{N+}Y_{N}\varphi_{N+}^{-1},\varphi_{\infty+}Y_{\infty}\varphi_{\infty+}^{-1}\bigr{)}.$
@@
The orbit of the coadjoint action of $\mathbb{G}_{R}=\!\bigl{(}\widetilde{G}_{D}^{(0)}\bigr{)}_{R}$ on an element ${\iota_{\lambda}\Lambda\in V_{-}=\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}}$ has the explicit form
$\mathcal{O}_{\Lambda}=\bigl{\{}\mathcal{P}_{-}\bigl{(}\operatorname{Ad}_{\varphi_{+}}^{\ast}\iota_{\lambda}\Lambda\bigr{)}\mid\varphi_{+}\in\widetilde{G}_{D+}^{(0)}\bigr{\}}.$
@@
By definition, the coadjoint orbit of $\bigl{(}\widetilde{G}_{D}^{(0)}\bigr{)}_{R}$ on any $\iota_{\lambda}\Lambda\in\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$ is given by
$\mathcal{O}_{\Lambda}=\bigl{\{}\operatorname{Ad}_{\varphi}^{R\ast}\iota_{\lambda}\Lambda\mid\varphi\in\bigl{(}\widetilde{G}_{D}^{(0)}\bigr{)}_{R}\bigr{\}}.$
Using the explicit form (2.5) for the coadjoint action of $\mathbb{G}_{R}$, we have
$\operatorname{Ad}^{R\ast}_{\varphi}\iota_{\lambda}\Lambda=\mathcal{P}_{-}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{+}}\iota_{\lambda}\Lambda\bigr{)}+\mathcal{P}_{+}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda\bigr{)}=\mathcal{P}_{-}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{+}}\iota_{\lambda}\Lambda\bigr{)},$
where the last equality follows from the fact that $\iota_{\lambda}\Lambda\in V_{-}=\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$, so that $\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda\in V_{-}$ also and hence $\mathcal{P}_{+}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{-}}\iota_{\lambda}\Lambda\bigr{)}=0$. ∎
@@
In the remaining sections, we will put this setup to use to describe the Lax matrix of the cyclotomic Gaudin model (1.9) as a point in a coadjoint orbit $\mathcal{O}_{\Lambda}$ for some suitable $\Lambda\in\Omega_{D^{\prime}}^{\Gamma}$ and then introduce the ingredients from point 3 in Section 2 to derive the associated Lax equations.
@@ 3.2 Lax matrix
Our algebraic setup covers the case of the cyclotomic Gaudin model with arbitrary multiplicities. However, as mentioned at the start of this section, from now on we will restrict to the case with simple poles at all $\omega^{k}\zeta_{r}$, $k\in\{0,\dots,T-1\}$, $r\in\{1,\dots,N\}$, and double poles at the origin and at infinity, since this is the setting required for our examples in Section 4. The discussion in the remaining sections is easily generalised to the case of arbitrary multiplicities.
To describe a coadjoint orbit $\mathcal{O}_{\Lambda}\in\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$ from Lemma 3.5 corresponding to the Lax matrix (1.9) of the cyclotomic Gaudin model, we fix a non-dynamical element $\Lambda\in\Omega_{D^{\prime}}^{\Gamma}$ with the same pole structure as (1.9), namely we introduce
$\Lambda(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{N}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{r}}{\lambda-\omega^{k}\zeta_{r}}+\Lambda_{\infty}\in\Omega_{D^{\prime}}^{\Gamma}.$
@@
$\mathcal{P}_{-}\Biggl{(}\varphi_{+}\iota_{\lambda}\Biggl{(}\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{N}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{r}}{\lambda-\omega^{k}\zeta_{r}}+\Lambda_{\infty}\Biggr{)}\varphi_{+}^{-1}\Biggr{)}$
$\qquad{}=\iota_{\lambda}\circ\pi_{\lambda}^{(1)}\Biggl{(}\frac{A_{0}^{(0)}}{\lambda}+\frac{A_{0}^{(1)}}{\lambda^{2}},\frac{1}{T}\frac{A_{1}}{\lambda-\zeta_{1}},\dots,\frac{1}{T}\frac{A_{N}}{\lambda-\zeta_{N}},A_{\infty}\Biggr{)}$
$\qquad{}=\iota_{\lambda}\Biggl{(}\frac{A_{0}^{(0)}}{\lambda}+\frac{A_{0}^{(1)}}{\lambda^{2}}+\frac{1}{T}\sum_{r=1}^{T-1}\sum_{k=0}^{T-1}\frac{\sigma^{k}A_{r}}{\lambda-\omega^{k}\zeta_{r}}+A_{\infty}\Biggr{)}$
$\qquad{}\equiv\iota_{\lambda}L(\lambda),$ (3.26)
@@
$A_{0}^{(0)}=\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1}+\bigl{[}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1}\bigr{]},$
$A_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1},$
$A_{r}=\phi_{r}\Lambda_{r}\phi_{r}^{-1},\qquad r\in\{1,\dots,N\},$
$A_{\infty}=\Lambda_{\infty},$ (3.27)
with $\phi_{r}^{(0)}$ denoted by $\phi_{r}$, for $r\in\{1,\dots,N\}$, to simplify notations. Notice that we have the relations $\sigma\phi_{0}^{(0)}=\phi_{0}^{(0)}$ and $\sigma\phi_{0}^{(1)}=\omega\phi_{0}^{(1)}$ for the field elements which ensure that $A_{0}^{(0)}\in\mathfrak{g}^{(0)}$ and $A_{0}^{(1)}\in\mathfrak{g}^{(-1)}$. This gives us the desired parameterisation of the cyclotomic Gaudin Lax matrix